OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 8x - y = -9 4x - 3y = -22 4(-8x - 9) - 3y = -22 4x - 3(-8x - 9) = -22 4(8x + 9) - 3y = -22 4x - 3(8x + 9) = -22 @ganeshie8

OpenStudy (ganpat):

solve for y in 1st equation ... i.e. put y on one side and rest on another

OpenStudy (ganpat):

and then, substitute value of y in another equation..

OpenStudy (anonymous):

8x - y = -9 Subtract 8x to both sides: -y = -8x - 9 Multiply all terms by -1: y = 8x + 9 Substitute 8x + 9 for y in the 2nd equation: 4x - 3(8x + 9) = -22 Can you solve for x? @iamabarbiegirl

OpenStudy (anonymous):

so the answer is D? @iGreen

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Oh, I didn't realize I gave the answer..but yes that's correct.

OpenStudy (anonymous):

lol thanks